OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. 23 B. 3 C. -9 D. -53

OpenStudy (anonymous):

Ok, so let us think about this one. You first evaluate the inside function which is g(x).

OpenStudy (anonymous):

g(-5) = -4(-5) + 3

OpenStudy (anonymous):

What is that?

OpenStudy (anonymous):

After you get the result of g(x), then you can take that and use it in f(x)

OpenStudy (solomonzelman):

This is an example \(\LARGE\color{blue}{ \bf f(x)=4x+2 }\) AND \(\LARGE\color{blue}{ \bf g(x)=3x-2 }\) Find fog when x=2? \(\LARGE\color{blue}{ \bf f(g)x))=4(3x-2)+2 }\) \(\LARGE\color{blue}{ \bf f(g)x))=12x-8+2 }\) \(\LARGE\color{blue}{ \bf f(g)x))=12x-6 }\) \(\LARGE\color{blue}{ \bf f(g)x))=12(2)-6 }\) \(\LARGE\color{blue}{ \bf f(g)x))=24-6 }\) \(\LARGE\color{blue}{ \bf f(g)x))=18 }\)

OpenStudy (anonymous):

I was just walking him through the problem @SolomonZelman, but showing him an example would help :)

OpenStudy (solomonzelman):

jtryon, you were a little off though. Nice try. Look at my example, and tell me if you see your err.

OpenStudy (anonymous):

f(g) not g(x)

OpenStudy (solomonzelman):

Yes, 1st plug in g(x) into the x of f(x), then simplify, and then plug in -5 for x.

OpenStudy (anonymous):

Yes, I see it now

OpenStudy (anonymous):

thanks everyone :) i have one more

OpenStudy (solomonzelman):

READ CAREFULLY. you put in for x, into g(x), what the f(x) is equal to.

OpenStudy (anonymous):

@jtryon

OpenStudy (solomonzelman):

ok, I see you don't want to try to do the problem. I am out.

OpenStudy (anonymous):

i do ido

OpenStudy (solomonzelman):

So do what I told you then...

OpenStudy (anonymous):

(-4(-5))+3= 23

OpenStudy (anonymous):

23 for each x at f(x) and it will be (-2(23))-7=-53 so D right?/